OpenStudy (anonymous):

Suppose there are 6 men sitting in a circular table. 4 of them have white shirts on and 2 have red shirts on. What is the probability that the men with a red shirt on do not sit together?

OpenStudy (anonymous):

3/5

OpenStudy (anonymous):

I'm not quite sure though, so let's wait for a few other people to answer. Don't close the question. :)

OpenStudy (anonymous):

How did you get that?

OpenStudy (anonymous):

well the chances of a guy with a red shirt sitting at the table are 1... one of them is definitely sitting a spot in the circular table. Now there are two seats on either side of him, and 5 people left over. the chances of a person with a red shirt sitting on his left is 1/5, and the chances of the person sitting on his right, so has a red shirt is 1/5... since it doesn't matter which side the other person sits in.. we can add the two probabilities together (1/5)+(1/5)=2/5. 2/5 is probability the people wear the red shirts will sit together, meaning that (1-(2/5)) equals the probability that they will not be sitting together. Thus, the answer is 3/5.

OpenStudy (anonymous):

However, like I said previously, I might incorrect, so I'd wait for verification.

OpenStudy (anonymous):

Ehh... sorry about the grammar mistakes, I was rushing while typing. Ooops.

OpenStudy (anonymous):

No problem. It sounds right! I'll just make sure someone else can make sure, but it does seem correct!

OpenStudy (anonymous):

Alright then :)

OpenStudy (raden):

agree u can also use "Probability of the complement of an Event" P(A) = 1-P(A') (with P(A') is probability even red always sit together)) P(A) = 1 - 4!*2/5! P(A) = 1 - 2/5 = 3/5

OpenStudy (anonymous):

Thank you guys!

OpenStudy (raden):

yw